# Transfer Learning of Surrogate Models: Integrating Domain Warping and Affine Transformations

### v1

@@ 5.2. Ablation study of transferring GPR on BBOB
Fig. 5 shows an ablation study on 2D BBOB functions, analyzing the impact of using only the beta CDF warping function, approximating an exponential transformation without rotation or translation. Results are compared to reproduced code from (Pan et al., 2024), which uses only affine transformations. The performance is visualized using box plots of raw SMAPE values obtained from 10 repetitions on 2D BBOB functions, with a transfer dataset of 20 samples. These plots compare the transferred GPR and a model trained solely on the transfer dataset. Since the target functions differ across the three settings, the box plot distributions highlight that, for most BBOB functions, optimizing only the beta CDF parameters in the transferred GPR outperforms jointly optimizing beta CDF and affine parameters and optimizing affine parameters alone. This highlights that optimizing additional parameters (“Full”) increases the complexity and difficulty of the problem. Moreover, the transferred model consistently outperforms the model trained from scratch when optimizing only the beta CDF parameters across nearly all BBOB functions. Fig. 21 (refer to the supplementary material (Authors, s)) presents the results for the 5D case, which exhibits a similar pattern to the observations described above.
The inclusion of rotation and translation in the transformations can introduce boundary effects, where parts of the function landscape initially outside the domain are mapped into it after transformation (Pan et al., 2024). We also examine the ”in-domain” setting, where only transfer data—mapped back to the original domain post-transformation—was used for training. This analysis, presented in Fig. 6, complements the previously examined scenario of random sampling directly from the target function. This figure demonstrates how model performance varies with different transfer dataset sizes and sampling strategies. For most functions, both “2D Transferred” and “2D Transferred (in the domain)” exhibit decreasing SMAPE values as the number of samples increases, indicating improved performance with more data. Notably, certain functions, such as F6, F10, F11, and even the more challenging ones like F23 and F24, display significant performance gaps between the two models, emphasizing the benefits of staying within the domain. In contrast, for functions such as F2 and F7, the performance of “2D Transferred” and “2D Transferred (in domain)” is nearly identical, suggesting that domain restriction has minimal impact in these cases. Although some functions perform better on “Transferred (in domain)” than “Transferred” under specific transfer dataset sizes, Fig. 22 (see supplementary material (Authors, s)) shows that in the 5D case, the difference is minimal, unlike in 2D cases.
@@ 5.3. Transferring GPR on real-world benchmark from automobile industry
As highlighted in (Pan et al., 2025), optimizing with affine transformations has proven effective for many real-world transfer learning applications. However, transferring knowledge between problem instances in the automobile industry presents persistent challenges that require further investigation. We choose this highly challenging benchmark to showcase and evaluate the effectiveness of our transfer learning approach. Fig. 7 (with full results available in Fig. 23, as detailed in the supplementary material (Authors, s)) illustrates the SMAPE trends for four GPR models across varying transfer dataset sizes, highlighting a subset of the experimental results. These include the original GPR model, a transferred GPR model that assumes an unknown affine transformation between the source and target problem instances (referred to as “Transferred (Affine only)” (Pan et al., 2024)), a transferred GPR model leveraging our proposed method (referred to as “Transferred (Full)”), and a model trained solely on the transfer dataset.
@@ 6. Conclusion
Experiments with BBOB functions demonstrate the effectiveness of the proposed method. With 20-sample transfer datasets, the transferred GPR outperforms models trained from scratch, particularly in 10D settings. However, with more samples, especially in 5D, scratch-trained models eventually surpass transferred models. The benefits of transfer learning are limited for highly complex functions like F16 and F21–F24, where the original GPR struggled with accurate approximation.
The proposed transfer learning method is also validated on a highly challenging real-world automotive task, demonstrating its effectiveness in low-data scenarios. However, substantial gaps between the source and target domains limit the transfer’s effectiveness, even with larger transfer datasets. This highlights the ongoing challenge of adapting the method to different problem instances within the automotive domain and the need for a deeper understanding of their relations.
Future research includes (1) extending this approach to other regression models, such as random forests; (2) integrating active learning into the transfer learning process to sample data points, which will maximally increase the performance; (3) investigating the trade-off between the expressivity of the transformation and the trainability of the transfer learning method.

### v2

@@ 5.2. Ablation study of transferring GPR on BBOB
Fig. 5 shows an ablation study on 2D BBOB functions, analyzing the impact of using only the beta CDF warping function, approximating an exponential transformation without rotation or translation. Results are compared to reproduced code from (Pan et al., 2024), which uses only affine transformations. The performance is visualized using box plots of raw SMAPE values obtained from 10 repetitions on 2D BBOB functions, with a transfer dataset of 20 samples. These plots compare the transferred GPR and a model trained solely on the transfer dataset. Since the target functions differ across the three settings, the box plot distributions highlight that, for most BBOB functions, optimizing only the beta CDF parameters in the transferred GPR outperforms jointly optimizing beta CDF and affine parameters and optimizing affine parameters alone. This highlights that optimizing additional parameters (“Full”) increases the complexity and difficulty of the problem. Moreover, the transferred model consistently outperforms the model trained from scratch when optimizing only the beta CDF parameters across nearly all BBOB functions. Fig. 21 (refer to the supplementary material (Pan et al., 2025b)) presents the results for the 5D case, which exhibits a similar pattern to the observations described above.
The inclusion of rotation and translation in the transformations can introduce boundary effects, where parts of the function landscape initially outside the domain are mapped into it after transformation (Pan et al., 2024). We also examine the ”in-domain” setting, where only transfer data—mapped back to the original domain post-transformation—was used for training. This analysis, presented in Fig. 6, complements the previously examined scenario of random sampling directly from the target function. This figure demonstrates how model performance varies with different transfer dataset sizes and sampling strategies. For most functions, both “2D Transferred” and “2D Transferred (in the domain)” exhibit decreasing SMAPE values as the number of samples increases, indicating improved performance with more data. Notably, certain functions, such as F6, F10, F11, and even the more challenging ones like F23 and F24, display significant performance gaps between the two models, emphasizing the benefits of staying within the domain. In contrast, for functions such as F2 and F7, the performance of “2D Transferred” and “2D Transferred (in domain)” is nearly identical, suggesting that domain restriction has minimal impact in these cases. Although some functions perform better on “Transferred (in domain)” than “Transferred” under specific transfer dataset sizes, Fig. 22 (see supplementary material (Pan et al., 2025b)) shows that in the 5D case, the difference is minimal, unlike in 2D cases.
@@ 5.3. Transferring GPR on real-world benchmark from automobile industry
As highlighted in (Pan et al., 2025a), optimizing with affine transformations has proven effective for many real-world transfer learning applications. However, transferring knowledge between problem instances in the automobile industry presents persistent challenges that require further investigation. We choose this highly challenging benchmark to showcase and evaluate the effectiveness of our transfer learning approach. Fig. 7 (with full results available in Fig. 23, as detailed in the supplementary material (Pan et al., 2025b)) illustrates the SMAPE trends for four GPR models across varying transfer dataset sizes, highlighting a subset of the experimental results. These include the original GPR model, a transferred GPR model that assumes an unknown affine transformation between the source and target problem instances (referred to as “Transferred (Affine only)” (Pan et al., 2024)), a transferred GPR model leveraging our proposed method (referred to as “Transferred (Full)”), and a model trained solely on the transfer dataset.
@@ 6. Conclusion
Experiments with BBOB functions demonstrate the effectiveness of the proposed method. With 20-sample transfer datasets, the transferred GPR outperforms models trained from scratch, particularly in 10D settings. However, with more samples, especially in 5D, scratch-trained models eventually surpass transferred models. The benefits of transfer learning are limited for highly complex functions like F16 and F21–F24, where the original GPR struggled with accurate approximation. The proposed transfer learning method is also validated on a highly challenging real-world automotive task, demonstrating its effectiveness in low-data scenarios. However, substantial gaps between the source and target domains limit the transfer’s effectiveness, even with larger transfer datasets. This highlights the ongoing challenge of adapting the method to different problem instances within the automotive domain and the need for a deeper understanding of their relations.
Future research includes (1) extending this approach to other regression models, such as random forests; (2) integrating active learning into the transfer learning process to sample data points, which will maximally increase the performance; (3) incremental learning can also be considered that each new sample can be used to fit the previously transferred model; (4) investigating the trade-off between the expressivity of the transformation and the trainability of the transfer learning method; (5) introduce a unified penalty framework that discourages any mapping outside the valid regions of both source and target distributions in high-dimensional or real-world settings; (6) by integrating our transferred surrogate into a standard surrogate‐assisted optimizer, we can analyze whether the transfer learning not only yields a more accurate surrogate but also accelerates the overall optimization process; (7) Kumaraswamy warping (Cowen-Rivers et al., 2022) may offer a faster, more flexible alternative to the beta CDF in practice and is worth investigating.